OpenStudy (anonymous):

How many solutions will the following system of linear equations have? –2x + 4y = –7 y=-1/2x+5

OpenStudy (dawnr):

this is easy

OpenStudy (dawnr):

do you know how to do a system of equations?

OpenStudy (dawnr):

-2x+7=-4y

OpenStudy (anonymous):

1 solution. It cannot have 2 because they do not curve at all. It cannot have 0 because they are not parallel. And it cannot have infinitely many because they are not the same line.

OpenStudy (dawnr):

2y=-x+10

OpenStudy (dawnr):

x=10-2y then you put x here---> -2x+7=-4y and solve it